# FCMR: Robust Evaluation of Financial Cross-Modal Multi-Hop Reasoning

### v2

@@ 1 Introduction
In this work, we propose Financial Cross-Modal Multi-Hop Reasoning (FCMR), a novel benchmark designed to address the limitations of existing datasets in cross-modal multi-hop reasoning. FCMR provides multiple-choice QA samples that test the integration of facts from text, tables, and charts. It consists of three levels of difficulty: Easy, Medium, and Hard. As in Figure 1, all instances in FCMR necessitate understanding three modalities to be answered correctly. In addition, problems at the Hard level explicitly demand cross-modal three-hop reasoning, making them more challenging (Figure 2). As FCMR is built using data sources from the financial domain, it is relatively free from the risk of data contamination.
@@ 3.2 Procedure of CMRGen
Lastly, we apply two stages of lexical and syntactic paraphrasing using GPT-4o to enhance the diversity of expressions in statements. Afterward, we conduct LLM-based filtering, using Claude 3.5 Sonnet, to ensure semantic accuracy. By employing separate LLMs in each step, we aim to mitigate model-oriented biases. For Hard-level instances, human experts optionally review them to further refine quality.
@@ 3.3 Multiple-Choice Design
Previous research on cross-modal multi-hop reasoning has often employed descriptive or short-answer formats, evaluated with metrics like F1 or Exact Match. These approaches might result in inaccurate evaluations by penalizing semantically appropriate answers that slightly deviate in form. To address this, we adopt a multiple-choice format with three statements. Recent work Pang et al. (2024) argues that single-choice question formats are more suitable for model evaluation than free-form answers, supporting our decision. In contrast to existing multimodal benchmarks Li et al. (2023); Yue et al. (2024); Ying et al. (2024); Zhang et al. (2024b); Liu et al. (2024b) that typically rely on a single correct answer, our setup allows for zero to three correct statements. This strategy enhances the complexity of the reasoning process needed to answer the problem accurately, requiring a more comprehensive synthesis of all provided statements. It also allows for precise evaluation of models’ cross-modal multi-hop reasoning capabilities.
@@ 3.4 Data Quality Control
To uphold high data quality, we implement multifaceted verification protocols. [^6] Refer to Appendix C for full details of our strategies. Specifically, we utilize Word Position Deviation (WPD) and Lexical Deviation (LD) metrics Liu and Soh (2022) to evaluate paraphrasing quality and compare these values with those from MRPC Dolan and Brockett (2005) and PAWS Zhang et al. (2019). The outcomes, presented in Table 6, confirm the superiority of our paraphrasing method. Furthermore, to verify that our dataset avoids the contamination issue identified in MMQA, we replicate the contamination experiment under the same conditions. As shown in Table 8, when the chart images are withheld, GPT-4o’s performance approximates random guessing, alleviating the risk of data contamination in FCMR. We further mitigate potential biases by equalizing modality order, statement types, and answer distributions. Figure 21 demonstrates that our benchmark is well-balanced across various perspectives. The final dataset consists of 757 Easy, 728 Medium, and 714 Hard instances, demand 100% cross-modal three-hop reasoning, setting FCMR apart from MMQA and its variants.
@@ 4.2 Main Results
Experimental results indicate that open-source models with fewer than 8B parameters continue to perform comparably to random selection for tasks at the Medium and Hard levels. However, for the Easy category, they demonstrate superior performance compared to similarly sized MLLMs. Surprisingly, even advanced MLLMs such as GPT-4o and Claude 3.5 Sonnet achieve performance gains in certain cases, suggesting that their visual interpretation capabilities are still not perfect. In Section E, we dive deeper into this phenomenon.
@@ 5.2 Stage-Based Analysis
Level Text Table Chart Total
@@ 6.1 Error Rate by Modality
Table 4 displays the numbers and proportions of statements Claude fails to interpret correctly, based on random 90 statements for each difficulty level. [^10] Errors unrelated to modality, such as misinterpreting conditions, are excluded from this analysis. At the Easy level, Claude often struggles with analyzing charts, performing noticeably worse compared to its handling of text and tables. This disparity indicates that MLLMs exhibit weaker capabilities in interpreting charts than in processing textual or tabular data (see Appendix E for details). However, as the difficulty level increases, errors in text and tables become more prominent. This shift is likely attributed to the increased complexity of reasoning chains, which raises the likelihood of errors in processing text and tables.
@@ B.5 Input Data Construction: Text, Tables, and Charts
B.6 Distractor Generation
B.7 Input Chart Code Generation
@@ C.1 Paraphrasing Quality
To evaluate the quality of Lexical-Syntax 2-Stage Paraphrasing, we employed the Word Position Deviation (WPD) and Lexical Deviation (LD) metrics proposed in Liu and Soh (2022). WPD assesses the syntactic diversity of paraphrased sentences, while LD evaluates lexical diversity. For an objective comparison, as shown in Table 6, we benchmarked the WPD and LD metrics of FCMR against prominent paraphrasing datasets such as MRPC Dolan and Brockett (2005) and PAWS Zhang et al. (2019), demonstrating the superior quality of our paraphrasing. Additionally, we validated semantic consistency using Claude 3.5 Sonnet to filter out samples where the paraphrased sentences were flagged as semantically altered. For hard statements with longer sentence lengths, we considered the potential for ambiguity introduced by paraphrasing. Consequently, all instances were manually reviewed, and sentences with ambiguous meanings were revised accordingly.
@@ D.1 Trend Assessment Error
Difficulty Total line bar scatter pie
Easy 75.43 74.89 78.60 71.01 84.31
Medium 50.82 52.70 50.00 49.79 -
Hard 30.39 39.22 29.20 23.44 -
Dataset: FCMR (Hard) Image? Accuracy
Random Selection - 12.28
GPT-4o 14.71
24.37
Table 8: Replication of experiments from Table 2 with FCMR. Despite the inherently challenging nature of the benchmark, GPT-4o’s performance drops to near random selection when charts are omitted, suggesting that FCMR is relatively robust against data contamination.
@@ Appendix E Chart Interpretability
Table 9: Performance of o1-like MLLMs on 100 randomly selected FCMR Hard-level samples. Due to the high cost of o1-like models, we limited our evaluation to these 100 samples.
@@
Figure 22: Reasons for Inference Failures by Model Across Fine-Grained Reasoning Stages
@@
Table 12: Detailed Statement Types by Difficulty and Modality Types. For the Easy level, all three answer statements are single-modal one-hop, while for the Medium level, all three statements are cross-modal two-hop. At the Hard level, all three statements consist of cross-modal three-hop. Even if each answer statement is a one-hop, the overall question remains a cross-modal three-hop QA. Specific examples can be found in Table 10

### v4

@@ 1 Introduction
In this work, we propose Financial Cross-Modal Multi-Hop Reasoning (FCMR), [^1] Our dataset is available at https://github.com/HYU-NLP/FCMR. a novel benchmark designed to address the limitations of existing datasets in cross-modal multi-hop reasoning. FCMR provides multiple-choice QA samples that test the integration of facts from text, tables, and charts. It consists of three levels of difficulty: Easy, Medium, and Hard. As in Figure 1, all instances in FCMR necessitate understanding three modalities to be answered correctly. In addition, problems at the Hard level explicitly demand cross-modal three-hop reasoning, making them more challenging (Figure 2). As FCMR is built using data sources from the financial domain, it is relatively free from the risk of data contamination.
@@ 3.2 Procedure of CMRGen
Lastly, we apply two stages of lexical and syntactic paraphrasing using GPT-4o to enhance the diversity of expressions in statements. Afterward, we conduct LLM-based filtering, using Claude 3.5 Sonnet, to ensure semantic accuracy. By employing separate LLMs in each step, we aim to mitigate model-oriented biases. Our framework also allows for an optional human expert review to improve their quality further. For FCMR, we use this review step to ensure high standards, especially for challenging (i.e., Hard-level) problems. Appendix C.1 provides the specific details of this human annotation process.
3.3 Distractor Generation
@@ 3.4 Multiple-Choice Design
Previous research on cross-modal multi-hop reasoning has often employed descriptive or short-answer formats, evaluated with metrics like F1 or Exact Match. These approaches might result in inaccurate evaluations by penalizing semantically appropriate answers that slightly deviate in form. To address this, we adopt a multiple-choice format with three statements. Recent work Pang et al. (2024) argues that single-choice question formats are more suitable for model evaluation than free-form answers, supporting our decision. In contrast to existing multimodal benchmarks Li et al. (2023); Yue et al. (2024); Ying et al. (2024); Zhang et al. (2024b); Liu et al. (2024b) that typically rely on a single correct answer, our setup allows for zero to three correct statements. This strategy enhances the complexity of the reasoning process needed to answer the problem accurately, requiring a more comprehensive synthesis of all provided statements. It also allows for precise evaluation of models’ cross-modal multi-hop reasoning abilities.
Dataset WPD LD
MRPC 0.12 0.42
PAWS 0.07 0.13
FCMR (Ours) 0.2 0.45
Dataset: FCMR (Hard) Image? Accuracy
Random Selection - 12.28
GPT-4o 14.71
24.37
Table 4: Replication of the experiments from Table 2 with FCMR. GPT-4o’s performance drops to near random selection when charts are omitted, suggesting that FCMR is relatively robust against data contamination.
@@ 3.5 Data Quality Control
To uphold high data quality, we implement multifaceted verification protocols. [^6] Refer to Appendix C for full details of our strategies. Specifically, we utilize Word Position Deviation (WPD) and Lexical Deviation (LD) metrics Liu and Soh (2022) to evaluate paraphrasing quality and compare these values with those from MRPC Dolan and Brockett (2005) and PAWS Zhang et al. (2019). The outcomes, presented in Table 3, confirm the superiority of our paraphrasing method. Furthermore, to verify that our dataset avoids the contamination issue identified in MMQA, we replicate the contamination experiment under the same conditions. As shown in Table 4, when the chart images are withheld, GPT-4o’s performance approximates random guessing, alleviating the risk of data contamination in FCMR. We further mitigate potential biases by equalizing modality order, statement types, and answer distributions. Figure 21 demonstrates that our benchmark is well-balanced across various perspectives. The final dataset includes 757 Easy, 728 Medium, and 714 Hard instances. Hard-level problems demand precise cross-modal three-hop reasoning. While statements in Easy and Medium questions involve simpler reasoning (e.g., single-modal one-hop for Easy), resolving each instance still necessitates processing information from all three provided modalities—text, table, and chart.
3.6 Chart Type Coverage
To ensure sufficient diversity in chart types, FCMR includes four commonly used formats in the financial domain: line, bar, scatter, and pie charts. These chart types reflect the typical visualizations found in real-world financial documents. A recent analysis Christensen et al. (2024) of 23,906 quantitative infographics extracted from all 10-K reports filed with the SEC between 2003 and 2019 shows that bar charts account for 55.0%, pie charts for 30.6%, and line charts for 12.7%. The chart types included in our FCMR dataset directly correspond to these prevalent quantitative categories, collectively covering approximately 98% of the most frequently used quantitative visualization types in actual 10-K filings. Details on the chart generation process are provided in Appendix B.6.
@@ 4.2 Main Results
Experimental results indicate that open-source models with fewer than 8B parameters continue to perform comparably to random selection for tasks at the Medium and Hard levels. However, for the Easy category, they demonstrate superior performance compared to similarly sized MLLMs. Surprisingly, even advanced MLLMs such as GPT-4o and Claude 3.5 Sonnet achieve performance gains in certain cases, suggesting that their visual interpretation capabilities are still not perfect. In Section 6.2, we dive deeper into this phenomenon.
@@ 5.2 Stage-Based Analysis
Level / Modality Text Table Chart Total
@@
Level / Type Line Bar Scatter Pie Total
Easy 74.89 78.60 71.01 84.31 75.43
Medium 52.70 50.00 49.79 - 50.82
Hard 39.22 29.20 23.44 - 30.39
@@ 6.1 Error Rate by Modality
Table 6 displays the numbers and proportions of statements Claude fails to interpret correctly, based on random 90 statements for each difficulty level. [^10] Errors unrelated to modality, such as misinterpreting conditions, are excluded from this analysis. At the Easy level, Claude often struggles with analyzing charts, performing noticeably worse compared to its handling of text and tables. This disparity indicates that MLLMs exhibit weaker capabilities in interpreting charts than in processing textual or tabular data (see Appendix 6.2 for details). However, as the difficulty level increases, errors in text and tables become more prominent. This shift is likely attributed to the increased complexity of reasoning chains, which raises the likelihood of errors in processing text and tables.
6.2 Chart Interpretability
@@ 7 Conclusion
Acknowledgments
This work was supported by the National Research Foundation of Korea(NRF) grant funded by the Korea government(MSIT) (RS-2025-00558151). This work was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) under the artificial intelligence semiconductor support program to nurture the best talents (IITP-(2025)-RS-2023-00253914) grant funded by the Korea government(MSIT). This work was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) (No.RS-2020-II201373, Artificial Intelligence Graduate School Program(Hanyang University)).
@@ B.5 Input Data Construction: Text, Tables, and Charts
B.6 Input Chart Code Generation
@@ C.1 Paraphrasing Quality
To evaluate the quality of Lexical-Syntax 2-Stage Paraphrasing, we employed the Word Position Deviation (WPD) and Lexical Deviation (LD) metrics proposed in Liu and Soh (2022). WPD assesses the syntactic diversity of paraphrased sentences, while LD evaluates lexical diversity. For an objective comparison, as shown in Table 3, we benchmarked the WPD and LD metrics of FCMR against prominent paraphrasing datasets such as MRPC Dolan and Brockett (2005) and PAWS Zhang et al. (2019), demonstrating the superior quality of our paraphrasing. Additionally, we validated semantic consistency using Claude 3.5 Sonnet to filter out samples where the paraphrased sentences were flagged as semantically altered.
To ensure high quality and robustness in FCMR’s Hard-level instances, a specific human annotation and review process was employed. Human reviewers initially examined 10% of samples from all difficulty levels (Easy, Medium, and Hard). This initial review showed that Easy and Medium samples were of high quality, so no further extensive human annotation was deemed necessary for them. Due to their complexity, Hard-level questions underwent a more rigorous human review to ensure their quality. Through this focused review, 22 out of 714 Hard-level samples (approximately 3%) were identified for revision. These revisions aimed to improve clarity, correctness, or the intended reasoning complexity. Finally, two authors cross-checked each revision to ensure consensus and maintain high-quality standards.
@@ D.7 Problem Understanding & Condition Ignoring
nova-pro-v1 Amazon (2024) 22
grok-2-vision xAI (2024) 19
qwen2.5-vl-72b-instruct Qwen et al. (2025) 19
gemma-3-27b-it DeepMind (2025) 19
Table 9: Performance of various advanced Multimodal Large Language Models (MLLMs), including o1-like reasoning models and other recent models, on 100 randomly selected FCMR Hard-level samples. Evaluation was limited to this 100-sample subset due to the operational costs associated with these models.
@@
Figure 22: Distribution of inference failures for each MLLM across the four reasoning stages (Planning, Modality Identification, Information Retrieval, and Information Reasoning), based on manual analysis of 40 Hard-level samples.
@@
Table 12: Detailed Statement Types by Difficulty and Modality Types. For the Easy level, all three answer statements are single-modal one-hop, while for the Medium level, all three statements are cross-modal two-hop. At the Hard level, all three statements consist of cross-modal three-hop. For Easy level questions, while each of the three statements is single-modal one-hop, determining the correct overall answer for the instance (i.e., selecting all true statements) requires processing information from all three provided modalities (text, table, and chart), ensuring comprehensive multi-modal engagement. Specific examples can be found in Table 10.